OpenStudy (anonymous):

Mulitply. Simplify your answer when possible -3x(x^3 + 2x^2 -x)

OpenStudy (anonymous):

I'm lost for some reason :(

OpenStudy (whpalmer4):

Okay, just use the distributive property here. \[a(b+c) = ab + ac\] I'll solve a similar problem, so as not to give you the answer: \[-2x(x^2 -x) = -2x*x^2 - 2x*(-x) = -2x^3 +2x^2\] Note well that change of sign in the final term! You have to be very careful when multiplying a negative something times something in parentheses — make sure you always apply the negative sign from the multiplier on the outside, and subtracting a negative number is the same as adding.

OpenStudy (anonymous):

ok thank you so so much I will study this for a bit to try and figure it out.